"please help me solve this problem. If a number n is added to 19, the result is 5 the number n. what is the number.
Answer by uday1435(57)

Let us write this statement into an equation
19 + n = 5
Subtracting 19 on both sides of this equation we get
19 + n – 19 = 5 – 19
So n = 5 – 19 = - 14
Answer: The number is – 14


If you still require further clarifications please write to udayakumar.t.r@gmail.com